# Empowering LLMs to Understand and Generate Complex Vector Graphics

### v2

@@ 5.2 Qualitative Evaluation
Figure 3 illustrates the visual quality of SVG generation methods, comparing both optimization-based and LLM-based approaches. It is evident that our method outperforms other LLM-based methods in terms of the completeness of the SVG generation, the selection and placement of primitives, and the semantic richness conveyed by the vector graphics. Optimization-based methods [11, 54, 66, 23, 67] use samples from the Latent Diffusion Model as supervision during the SVG generation process. Consequently, these methods normally employ a large number of overlapping and interwoven primitives to closely approximate the realistic samples. This often lead to excessive stroke redundancy, and the individual shapes of the primitives may appear irregular when viewed in isolation, making them less practical for use in real-life applications. In Fig. S7, we show more SVGs generated by our LLM4SVG.
@@
Human Evaluation. To evaluate the visual effects of our generated SVGs compared to other LLMs, we conducted a user study. As Claude 3.5 and GPT-4 are currently recognized as two of the best LLMs, we compared SVGs generated by our LLM4SVG with those produced by these models. In this study, we shuffled the SVGs generated by the different models together with the selected examples from our dataset. Participants were only provided with the text descriptions and the shuffled SVGs. They were then asked to evaluate the SVGs based on prompt alignment, visual quality, and their willingness to use the SVGs in real-world scenarios. The results are presented in Table 3, where our generated SVGs significantly outperformed other LLM-based methods, and the visual quality of our SVGs is closely comparable to that of SVGs created by humans. Further details of human evaluation are provided in Sec. C of Supplementary.
Figure 4: Our LLM4SVG Model Generates SVGs with Primitive Ordering that Aligns with Human Design Principles. The upper example demonstrates a gradual design process, progressing from individual components to the complete design. In contrast, the lower example begins with the overall design before detailing each individual component.
@@ 5.3 Ablation Study
Analysis of the LLM4SVG Architecture. The performance of LLM4SVG is significantly influenced by the underlying base model, especially concerning how the LLM tokenizer manages continuous numeric tokens. For example, in the Qwen models [59, 52], during the tokenization process, all numbers and decimal points are converted into a single token. As a result, Qwen is unable to properly handle continuous numerical data, leading to poor performance in tasks such as coordinate and hexadecimal color prediction, making it difficult to generate complete and coherent SVGs. LLMs like GPT-2 [39], which enumerate all numbers within 1,000 as individual tokens in the tokenizer, exhibit a stronger understanding of numerical data. This capability allows such models to generate more complete and coherent SVGs, leading to a more visually appealing and harmonious output.
Analysis of SVG Semantic Tokens. During the model training process, we expanded the tokenizer by incorporating 55 SVG-specific semantic tokens. This modification enables SVG tags and attributes to be identified with distinct tokens, thereby differentiating them from general text. For instance, prior to this addition, the word “path” could ambiguously refer to an SVG tag or merely a physical path. However, with the introduction of a dedicated path token, the model is now equipped to clearly differentiate between the SVG path tag and a literal path. This distinction aids the model in more effectively and accurately learning the SVG generation template, enhancing its ability to produce precise and contextually appropriate SVGs.
Figure 5: t-SNE Visualization of Token Embeddings. The green dots represent the SVG token embeddings initialized with descriptions, while the yellow crosses indicate those initialized without descriptions. The blue squares represent SVG token embeddings after training.
Analysis of Word Embedding Initialization Method. As illustrated in Fig. 5, we used T-SNE [57] to map the newly added tokens into a two-dimensional visualization for better demonstration. In this visualization, the green dots represent tokens that were initialized using the text description of each token. Specifically, we averaged the values obtained from tokenizing these descriptions to initialize the tokens, a process detailed in Sec. 4.1. This strategy groups the token embeddings into a relatively compact region within the feature space, which helps reduce the difficulty of model training. The yellow crosses, on the other hand, indicate tokens that were initialized without using the text description. These tokens exhibits a more scattered distribution across the feature space, making it challenging for the model to learn the accurate meaning of these tokens. The blue squares represent the positions of the tokens within the feature space post-training. It is evident that after the training phase, the token embeddings not only maintain local continuity, evidenced by smooth transitions within clusters, but also exhibit a reasonable of discreteness, where different functional SVG elements form distinct clusters. Additionally, these tokens remain closely within the overall feature space. This visualization demonstrates the rationale behind our token addition method and the effectiveness of our training approach.
@@ 6 Conclusion & Discussion
In this work, we proposed a method that harnesses the capabilities of Large Language Models (LLMs) for both understanding and generating SVGs, enabling LLMs to directly interpret SVG source code and generate high-quality SVGs. These generated SVGs exhibit a certain level of complexity, and align with human design principles. Our method introduces a structured SVG encoding approach that addresses the challenge of LLMs treating SVG source code as ordinary text. Traditionally, this treatment often results in inefficient use of tokens for learning SVG structures and numerical encoding, thereby complicating model training. Additionally, we have collected a dataset consisting of 250,000 high-quality SVGs designed by human designers and 580,000 SVG-text instruction pairs. This dataset is expected to significantly expedite the future research related to vector graphics.
Future Work. Our LLM4SVG can leverage different LLMs as its base model. In the future, more advanced and complex models could serve as the base, and combined with multimodal large language model to integrate image information. This extension may help reduce the difficulty of model understanding and generation, enabling more accurate and efficient SVG generation.
@@
In this supplementary material, we provide additional details and discussions related to our work on LLM4SVG. Specifically, it covers the following aspects:
  • In Section A, we demonstrate the advantages of our approach over existing LLM-based methods for SVG generation, particularly in terms of visual appeal and the ability to handle numerical coordinates.
  • In Section B, we show examples from our newly collected dataset and introduce a lossless preprocessing pipeline.
  • In Section C, we provide more details about the user study, including the number of participants, their backgrounds, and how the metrics were obtained.
  • In Section D, we provide the details of the SVG semantic tokens.
  • In Section E, we present additional qualitative results of our LLM4SVG, showcasing its ability to generate SVGs with high editability and superior visual quality.
@@ Appendix A Comparison with LLM-based Methods
Additionally, we compare the SVG source codes generated by our method and two of the most recent LLMs, GPT-o1-preview [2] and Claude 3.5-sonnet [35], to demonstrate the superiority of our method. As shown in Fig.S4, given the prompt “umbrella”, both GPT-o1-preview and Claude 3.5-sonnet can only predict integer coordinates, whereas our LLM4SVG is capable of generating precise decimal coordinates accurate to two decimal places. In the context of SVG representation, retaining only integer values can lead to incomplete or distorted SVG shapes, as illustrated in Fig.S4. We present additional examples in Fig. S1 to further demonstrate the visual differences between integer and decimal coordinates.
@@
Figure S2: Samples from Our Collected SVG Dataset. Our collected datasets are extensive and highly diverse.
Figure S3: Qualitative Comparison of LLM4SVG with Existing LLM-Based Methods.
Figure S4: Example SVG Code Comparison. SVG Code Generated by GPT-o1-preview [2], Claude 3.5-sonnet [35], and Our LLM4SVG.
Figure S5: Illustration of our SVG Processing Pipeline.
Appendix B Dataset and Preprocessing Pipeline
In Fig. S2, we present examples from our extensive and diverse SVG dataset. This dataset includes primitives with varying levels of complexity, ranging from minimal to highly detailed, and utilizes a rich and harmonious color palette. Furthermore, it encompasses a wide range of subjects, including people, animals, objects, and symbols.
As discussed in Sec. 3 of our manuscript, approximately half of the data in an SVG file is redundant for visual rendering. Consequently, we introduce a lossless SVG preprocessing pipeline to streamline the dataset and enhance training efficiency. As illustrated in Fig. S5, we remove redundant elements of SVGs, such as XML declarations, comments, titles, descriptions, <def> and <class> tags, and global <g> tags. We also resize the canvas size to $128\times 128$, along with adjusting all coordinates of each primitive. Finally, we round decimals to a maximum of two decimal places. These operations preserve the visual output of the SVGs while improving the efficiency of training.
Following the preprocessing of the SVGs, we structured the training instruction template for both SVG generation and understanding tasks. For understanding tasks, we converted all SVGs into raster images and employed GPT-4 [2] to generate detailed descriptions, which serve as the learning content. For generation tasks, we utilized captions generated by BLIP [26] as prompts, with the SVG source code serving as the learning content. In total, our dataset comprises 580k unique high-quality SVG-text instruction-compliant samples.
Figure S6: Screenshot of Our Questionnaires.
@@ Appendix B Dataset and Preprocessing Pipeline
SVG Shape Tags [<|svg_path|>] svg path element
@@
Figure S7: Additional Results Generated by Our LLM4SVG.
Appendix C More Details about User Study
@@
Appendix D SVG Semantic Tokens
Appendix E Additional Qualitative Results
In Fig. S7, we present additional results generated by LLM4SVG. These examples demonstrate the model’s capability to produce SVGs that not only maintain semantic integrity but also exhibit high visual quality.

### v3

@@ 5.3 Qualitative Evaluation
Figure 4 illustrates the visual quality of SVG generation methods, comparing both optimization-based and LLM-based approaches. It is evident that our method outperforms other LLM-based methods in terms of the completeness of the SVG generation, the selection and placement of primitives, and the semantic richness conveyed by the vector graphics. Optimization-based methods [14, 53, 64, 25, 65] use samples from the Latent Diffusion Model as supervision during the SVG generation process. Consequently, these methods normally employ a large number of overlapping and interwoven primitives to closely approximate the realistic samples. This often lead to excessive stroke redundancy, and the individual shapes of the primitives may appear irregular when viewed in isolation, making them less practical for use in real-life applications.
@@
Human Evaluation. To evaluate the visual effects of our generated SVGs compared to other LLMs, we conducted a user study. As Claude 3.5 and GPT-4 are currently recognized as two of the best LLMs, we compared SVGs generated by our LLM4SVG with those produced by these models. In this study, we shuffled the SVGs generated by the different models together with the selected examples from our dataset. Participants were only provided with the text descriptions and the shuffled SVGs. They were then asked to evaluate the SVGs based on prompt alignment, visual quality, and their willingness to use the SVGs in real-world scenarios. The results are presented in Table 3, where our generated SVGs significantly outperformed other LLM-based methods, and the visual quality of our SVGs is closely comparable to that of SVGs created by humans. Further details of human evaluation are provided in Sec. D of Supplementary.
@@ 5.4 Ablation Study
Analysis of the LLM4SVG Architecture. The performance of LLM4SVG is significantly influenced by the underlying base model, particularly in how the tokenizer processes continuous numeric tokens. For instance, in the Qwen models [57, 51], all numbers and decimal points are tokenized as a single unit. This approach prevents Qwen from effectively handling continuous numerical data, resulting in poor performance in tasks such as coordinate prediction and hexadecimal color generation. Consequently, the model struggles to produce complete and coherent SVGs. In contrast, LLMs like GPT-2 [40], which explicitly enumerate all numbers up to 1,000 as individual tokens, demonstrate a stronger numerical understanding. This design enables them to generate more precise coordinates and color values, leading to more visually coherent and aesthetically refined SVGs.
Analysis of SVG Semantic Tokens. To enhance SVG understanding, we expanded the tokenizer with 55 SVG-specific semantic tokens, allowing the model to distinguish SVG tags and attributes from general text. For instance, without this modification, the word “path” could ambiguously refer to an SVG tag or a physical pathway. With a dedicated <path> token, the model can now differentiate between the two, improving SVG generation accuracy.
We analyze the impact of these tokens using t-SNE visualization in Sec. C of Supplementary. The green dots represent embeddings initialized with descriptions, forming structured clusters for SVG geometry and path command tags. In contrast, the orange crosses, which represent randomly initialized tokens, are scattered, leading to slower convergence. After training, as shown by the blue squares, the embeddings become well-organized, demonstrating that the model effectively learns semantic relationships between SVG elements.
@@ 6 Conclusion & Discussion
In this work, we introduced LLM4SVG, the first framework designed to support existing LLMs/MLLMs in both understanding and generating SVGs. Our approach allows language models to directly interpret SVG source code and produce high-quality SVGs that demonstrate meaningful complexity and align with human design principles. Our methodology employs a structured SVG encoding strategy that overcomes the limitations of LLMs treating SVG source code merely as plain text. Additionally, we developed the SVGX-SFT Dataset, which comprises high-quality SVGs created by human designers and dialogue-based instruction pairs tailored specifically for training LLMs. We anticipate that this dataset will significantly accelerate future research in vector graphics.
Future Work. LLM4SVG lays the foundation for language models in vector graphics, with several promising directions for future exploration. Enhancing semantic understanding of SVG elements could enable more meaningful editing and generation. Bridging text, raster images, and vector graphics through cross-modal translation would improve creative workflows. Supporting iterative refinement via natural language feedback could make SVG generation more interactive and user-friendly. Additionally, adapting the approach to specific domains like data visualization or UI/UX design could enhance its practical applications.
Acknowledgment. This work was supported in part by the National Science and Technology Major Project (No. 2022ZD0117800), the Young Elite Scientists Sponsorship Program by the Chinese Association for Science and Technology (CAST), and the Fundamental Research Funds for the Central Universities. It was also partially funded by the National Natural Science Foundation of China (No. 62461160331, No. 62132001), the Hong Kong Research Grants Council (RGC) General Research Fund (No. 17203023), and the Hong Kong RGC Collaborative Research Fund (No. C5052-23G). Additionally, this research received support from the National Natural Science Foundation of China/RGC Collaborative Research Scheme (No. CRS_HKU703/24), the Huawei-BUAA Joint Lab, and the Hong Kong Jockey Club Charities Trust under Grant No. 2022-0174.
@@
In this supplementary material, we provide additional details and discussions related to our work on LLM4SVG. Specifically, this document covers the following aspects:
  • Comparison with Existing LLM-Based Methods (Sec. A): We demonstrate the advantages of our approach over existing LLM-based methods for SVG generation, particularly in terms of visual appeal and the ability to handle numerical coordinates.
  • Dataset and Preprocessing Pipeline (Sec. B): We introduce our newly collected dataset and describe a lossless preprocessing pipeline that enhances data quality for training and evaluation.
  • SVG Semantic Tokens (Sec. C): We provide details on the SVG semantic tokens used in our model, along with an analysis of how different word embedding initialization methods affect model performance.
  • User Study Details (Sec. D): We present additional details about the user study, including the number of participants, their backgrounds, and the methodology used to obtain evaluation metrics.
  • Primitive Ordering in SVG Generation (Sec. E): We illustrate how our LLM4SVG model generates SVGs by following a structured primitive ordering strategy. We compare two different generation approaches—a step-by-step composition process and a top-down refinement method—and discuss their implications for SVG interpretability and usability.
  • Additional SVG Generation Results (Sec. F): We showcase a diverse set of SVG illustrations generated by our model, demonstrating its ability to produce high-quality, semantically accurate, and stylistically consistent vector graphics across various categories, including objects, animals, food, and abstract symbols.
@@ Appendix A Comparison with LLM-based Methods
Additionally, we compare the SVG source codes generated by our method and two of the most recent LLMs, GPT-o1-preview [2] and Claude 3.5-sonnet [4], to demonstrate the superiority of our method. As shown in Fig. S4, given the prompt “umbrella”, both GPT-o1-preview [2] and Claude 3.5-sonnet [4] can only predict integer coordinates, whereas our LLM4SVG is capable of generating precise decimal coordinates accurate to two decimal places. In the context of SVG representation, retaining only integer values can lead to incomplete or distorted SVG shapes, as illustrated in Fig. S4. We present additional examples in Fig. S1 to further demonstrate the visual differences between integer and decimal coordinates.
@@
Figure S2: Illustrative Samples from the SVGX-SFT Dataset, showcasing its diversity in style, structure, and semantics. The dataset includes a wide range of objects, icons, and illustrations, making it well-suited for training and fine-tuning vector graphic generation models.
Figure S3: Qualitative Comparison of LLM4SVG with Existing LLM-based Methods. Given textual prompts (left), various LLMs generate corresponding vector graphics. The comparison highlights differences in abstraction, structural consistency, and fidelity to the input descriptions. Our LLM4SVG demonstrates improved coherence, accuracy, and stylistic refinement in SVG generation.
Figure S4: Example SVG Code Comparison For the Prompt “umbrella”. The figure contrasts SVG outputs generated by GPT-4o-preview, Claude 3.5-sonnet, and our LLM4SVG. While GPT-4o-preview and Claude 3.5-pro produce basic umbrella-like shapes, their structures contain artifacts or lack refinement. In contrast, LLM4SVG generates a more polished, visually coherent, and structured SVG representation, demonstrating improved detail, smoothness, and geometric accuracy.
Figure S5: Illustration of our SVG Processing Pipeline. The left side shows the original SVG code, which contains redundant elements such as XML declarations, comments, metadata, unused tags, and absolute coordinates. The right side presents the optimized SVG code after applying our cleaning pipeline, which improves efficiency, readability, and scalability by removing unnecessary elements, converting absolute coordinates to relative, and standardizing canvas size. The rendered output remains visually consistent while significantly reducing file complexity.
Appendix B SVGX-SFT Dataset Details
Figure S2 presents examples from our extensive and diverse SVGX-SFT dataset. This dataset includes primitives of varying complexity, ranging from minimal to highly detailed, while maintaining a rich and harmonious color palette. Additionally, it covers a broad spectrum of subjects, including people, animals, objects, and symbols, making it a comprehensive resource for both SVG generation and understanding tasks.
Preprocessing Pipeline. As discussed in Section 3 of our manuscript, a significant portion of an SVG file consists of redundant metadata that does not contribute to its rendered appearance. To improve training efficiency while preserving visual fidelity, we introduce a lossless SVG preprocessing pipeline, as illustrated in Figure S5.
Our pipeline systematically removes unnecessary elements from SVG files, including XML declarations, comments, titles, descriptions, <defs> and <class> tags, as well as global <g> tags. Additionally, we optimize numerical representations by converting absolute coordinates to relative ones and rounding decimal values to a maximum of two decimal places. Furthermore, all SVGs are resized to a standardized $128\times 128$ canvas, ensuring consistency across the dataset. These optimizations significantly reduce file size and computational overhead without altering the visual output.
Instruction-Based Dataset Construction. After preprocessing, we structured the dataset to support both SVG generation and understanding tasks. For understanding tasks, each SVG was rasterized into an image, and GPT-4 [2] was used to generate detailed descriptions, which serve as learning targets. For generation tasks, textual prompts were generated using BLIP [29], with the corresponding SVG source code serving as the learning content.
In total, our dataset comprises 580k high-quality SVG-text instruction-compliant samples, providing a robust foundation for training models in structured vector graphic generation and interpretation.
@@ Appendix B SVGX-SFT Dataset Details
SVG Geometry Tags [<|svg_path|>] svg path element
@@
Appendix C Our Proposed SVG Semantic Tokens
Figure S6: t-SNE Visualization of Token Embeddings. The green dots represent the SVG token embeddings initialized with descriptions, while the orange crosses indicate those initialized without descriptions. The blue squares represent SVG token embeddings after training.
Analysis of Word Embedding Initialization Method. As illustrated in Fig. S6, we used T-SNE [55] to map the newly added tokens into a two-dimensional visualization for better demonstration. In this visualization, the green dots represent tokens that were initialized using the text description of each token. Specifically, we averaged the values obtained from tokenizing these descriptions to initialize the tokens, a process detailed in Sec. 4.1. This strategy groups the token embeddings into a relatively compact region within the feature space, which helps reduce the difficulty of model training. The orange crosses, on the other hand, indicate tokens that were initialized without using the text description. These tokens exhibit a more scattered distribution across the feature space, making it challenging for the model to learn the accurate meaning of these tokens. The blue squares represent the positions of the tokens within the feature space post-training. It is evident that after the training phase, the token embeddings show more meaningful groupings, where tokens with similar semantics are clustered together while maintaining relative proximity to all SVG tokens. Additionally, these tokens remain closely within the overall feature space. This visualization demonstrates the rationale behind our token addition method and the effectiveness of our training approach.
Figure S7: Screenshot of Our Questionnaire used in the LLM4SVG User Study. Participants evaluate five SVGs generated from a given text prompt based on three key criteria: (1) Prompt Alignment—how well the SVG matches the given prompt, (2) Visual Quality—the aesthetic appeal of the SVG, and (3) Pick Score—whether the participant would consider using the SVG in real-world applications. Ratings are provided on a 5-point scale, with an additional Yes/No selection for practical usability.
Appendix D More Details about User Study
@@
Appendix E Primitive Ordering in SVG Generation
Figure S8 illustrates how our LLM4SVG model generates SVGs by following a primitive ordering strategy that aligns with human design principles. This structured approach ensures that vector graphics are constructed in an intuitive and interpretable manner.
The upper sequence in Figure S8 demonstrates a step-by-step composition process. The design begins with a few basic primitives, such as lines and simple shapes, and progressively adds more details. This method closely resembles how human designers create illustrations—starting with foundational elements before refining them into a complete object. For example, the umbrella starts with a simple handle, then adds canopy sections until the final structured design emerges. Similarly, the dolphin illustration begins with a basic shape, followed by gradual refinements to enhance realism. This process ensures that each step maintains logical continuity, making the SVG more comprehensible and easier to manipulate.
In contrast, the lower sequence showcases an alternative approach where the entire object is introduced first, followed by successive refinements. This method mimics a top-down design strategy, where an overall form is quickly established before adding intricate details. While this approach can be efficient for certain applications, it lacks the structured progression seen in human sketching workflows, which typically emphasize incremental construction.
By structuring SVG generation in a way that mirrors human cognitive processes, our LLM4SVG model enhances the interpretability and usability of vector graphics. This structured ordering makes the model particularly useful for applications such as educational tools that teach step-by-step drawing, design software that supports intuitive vector editing, and automated graphic generation systems that produce clear and logically constructed icons. The ability to generate illustrations progressively ensures that the output is both aesthetically pleasing and functionally adaptable.
Figure S8: Our LLM4SVG model generates SVGs with primitive ordering that aligns with human design principles. The upper example demonstrates a gradual design process, progressing from individual components to the complete design. In contrast, the lower example begins with the overall design before detailing each individual component.
Appendix F Additional Results Generated by Our LLM4SVG
Figure S9 showcases a diverse collection of SVG illustrations generated by our LLM4SVG model. These results demonstrate the model’s capability to produce high-quality, semantically accurate, and visually appealing vector graphics across a wide range of categories. The generated SVGs span various domains, including:
  • Objects: Everyday items such as a pen, paperclip, key, and teapot.
  • Animals: Illustrations of a dolphin, parrot, horse, fish, and giraffe.
  • Food: Fruits, vegetables, and prepared dishes, including a tomato, orange, lime, and a rice bowl.
  • People and Expressions: Human figures representing different professions, emotions, and activities.
  • Symbols and Abstract Concepts: Medals, graphs, weather symbols, and a heart icon.
These results highlight the effectiveness of our model in generating stylistically consistent and visually coherent vector graphics. The illustrations maintain a balanced level of abstraction while preserving key details, making them suitable for real-world applications such as digital icons, UI elements, and educational materials.
Figure S9: Additional results generated by our LLM4SVG model. This collection showcases a diverse range of high-quality SVG illustrations covering various categories, including objects, animals, food, people, symbols, and abstract concepts. The results demonstrate the model’s ability to produce visually appealing, semantically accurate, and stylistically consistent vector graphics.